Attached is the latest full noise budget we have. I generated it by rerunning current noise budget notebook. The aim of this post is to create a checkpoint of where we are, how the analysis was done and what we are going to do in the future.
There are 5 measurements of ASD of the beat note currently. 1 set (dark red) is from un-dated data (committed as ported from 40mSVN). Others have dates marked on them. These data sets are the spectrum of beat note frequency. No post-processing is done on these measurements and they are plotted as measured.
Coating Brownian Noise (Green):
Coating Thermo-Optic Noise (Pink):
: Profile-weighted temperature fluctuation PSD, calculated using:
: Effective thermoelastic coefficient for coating, calculated using Evans et a. (2008) Eq A1 and A4:
Here, : Thermoelastic coefficient for the k-th layer. and
: Thickness of each layer. This is read from coatingLayers.csv which contains optimized. There are 57 layers in total.
Rest variables are the same as previously stated.
: Effective thermo-refractive coefficient for coating, calculated using Evans et a. (2008) Eq B3-B8. This is bit tedious calculations so I'll leave out the details.
: Wavelength of beam light = 1064 nm
: Effective heat capacity per unit volume of the coating, where is the heat capacity of the kth layer.
: Heat capacity per unit volume of substrate =
Rest variables are as stated before.
Substrate Brownian Noise (Yellow):
Substrate Thermo Elastic Noise (Sky Blue):
PDH Shot Noise (Orange):
PLL Oscillation Noise (Grey):
PLL Readout Noise (Purple):
Seismic Noise (Black):
Photothermal Noise (Brown):
Residual NPRO Noise (Green):
We should be able to finish aligning South Path with the newly installed PMC by end of this week. Then if everything works fine, we should start taking beat note measurements every week and focus on updating old data in the estimate plots.